OpenStudy (anonymous):

need help.. T.T Given a+b=2c, evaluate: [2^((a-b)b) * 2^((b-c)(c-a) all over 2^((c+b)(c-b))] ^(1/c)

OpenStudy (freckles):

\[(\frac{2^{(a-b)b} \cdot 2^{(b-c)(c-a)}}{2^{(c+b)(c-b)}})^{^\frac{1}{c}}\]

OpenStudy (freckles):

is that one part write?

OpenStudy (anonymous):

yes thats it

OpenStudy (freckles):

are we suppose to write it in terms of just c?

OpenStudy (anonymous):

I dont really know.. T.T actually their are choices of the answer. a. 0 b. 1 c. 2 d. \[2^{2c}\]

OpenStudy (freckles):

Ok let's try something first use law of exponents and then also replace all the c's with (a+b)/2

OpenStudy (freckles):

I mean use the law of exponents on the product part the division part write just as 2^(some power) inside the parenthesis

OpenStudy (freckles):

I'm telling you to replace all the c's with (a+b)/2 because of the given part which is 2c=a+b

OpenStudy (freckles):

so first do you know what the properties say about simplifying \[2^m * 2^n\]

OpenStudy (anonymous):

\[2^{m+n}\]

OpenStudy (freckles):

ok and what if we had \[\frac{2^m \cdot 2^n}{2^c}\]

OpenStudy (anonymous):

\[2^{m+n-c}\]

OpenStudy (freckles):

So you would agree that we could write what you have as: \[(2^{(a-b)b +(b-c)(c-a)-(c+b)(c-b)})^\frac{1}{c}\]

OpenStudy (freckles):

took top exponents and add and subtracted the bottom exponent

OpenStudy (freckles):

I would multiply everything out and combine any like terms looking at \[(a-b)b+(b-c)(c-a)-(c+b)(c-b)\] Could you do that and let me know what you get ?

OpenStudy (anonymous):

\[bc+ac-2c\]

OpenStudy (freckles):

I think you might be missing some terms

OpenStudy (freckles):

so (a-b)b=ab-b^2 (b-c)(c-a)=bc-ab-c^2+ca (c+b)(c-b)=c^2-b^2

OpenStudy (freckles):

so you need to simplify \[ab-b^2+bc-ab-c^2+ca-(c^2-b^2)\] Try simplifying this

OpenStudy (freckles):

oh i see what you did you left off the square on the c so you meant to say bc+ac-2c^2

OpenStudy (anonymous):

ab-ab-b^2 +b^2 +bc +ca -c^2 -c^2 so only bc+ca-2c^2 remain

OpenStudy (freckles):

ok coolness! :)

OpenStudy (freckles):

Now let't replace all the c's in \[(2^{bc+ca-2c^2})^{^\frac{1}{c}}\] with (a+b)/2

OpenStudy (freckles):

so that means we now have: \[(2^{b \frac{a+b}{2}+\frac{a+b}{2}a-2(\frac{a+b}{2})^2})^{^\frac{2}{a+b}}\]

OpenStudy (freckles):

so we have one law of exponent that says we can do \[(2^{w+r+s})^v=2^{vw+vr+vs}\]

OpenStudy (freckles):

so that means we have \[2^{b \frac{a+b}{2} \frac{2}{a+b}+\frac{a+b}{2}a \frac{2}{a+b}-2(\frac{a+b}{2} )^2 \frac{2}{a+b}}\]

OpenStudy (freckles):

you should notice a lot of things cancel

OpenStudy (anonymous):

0 is the answer

OpenStudy (freckles):

how did you get that?

OpenStudy (freckles):

is that what the exponent simplified too?

OpenStudy (anonymous):

\[b+a-2(\frac{ a+b }{ 2 })\]

OpenStudy (anonymous):

b+a-a-b=0 :)

OpenStudy (freckles):

so you would have 2^0 which isn't 0 but ... hint: a^0=1 (well whenever we don't have a is 0)

OpenStudy (anonymous):

ahhh so it should be 1 thank you very much now i understand..:)

OpenStudy (freckles):

greatness yes the answer is 1